Circular No. 6691

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1997ct IN ANONYMOUS GALAXY S. Jha, P. Garnavich, P. Challis and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report that CCD images of SN 1997ct (cf. IAUC 6690) taken by M. Gomez on July 1.2 UT at the Whipple Observatory 1.2-m telescope yield V = 18.6 and precise position α = 14^{h}07^{m}26^{s}.88, δ = +70^{o}26′05′′.9 (equinox 2000.0, uncertainty ± 0′′.4), corresponding to an offset of 9′′.4 east and 1′′.4 south from the host galaxy’s center. A spectrogram taken by D. Koranyi with the 1.5-m telescope indicates that the supernova is of type Ia, four to six weeks after maximum. The host galaxy is at redshift 0.030, based on the observed Hα emission.

GRB 970616 M. A. Pahre, J. C. Clemens, S. R. Kulkarni, and I. N. Reid, Palo-

mar Observatory, California Institute of Technology; D. A. Frail, National Radio Astronomy Observatory; and J. Cordes, Cornell University, report: “Photographic IV-N plates acquired by J. Mueller and D. Mendenhall on the Palomar 1.2-m Oschin telescope (June 17.5 and 20.5 UT) of the optical transient (OT) identified by Galama et al. (IAUC 6687) in the field of GRB 970616 show no evidence for the OT to the plate limit of Cousins I ∼ 18. Palomar 1.52-m telescope observations of the OT (with a CCD and a nearinfrared camera) yield the following magnitudes: June 20.5, J = 18.30 ± 0.05, K = 16.9 ± 0.4; June 21.5, J = 18.23 ± 0.11; June 23.5, Cousins I ≥ 21.5 (3σ limit, at which it was marginally detected). Palomar 5-m Hale telescope observations of the OT on June 28.5 show the source to have J = 18.31 ± 0.04 and K = 17.06 ± 0.17. The source exhibited no detected variability in either the J or K bands between June 20 and 28, consistent with the nondetection of variability reported by Castro-Tirado et al. (IAUC 6688) in the K band. Thus, on June 21 the colors of the OT were R − J = +4.25 and J − K = +1; on June 28, J − K = +1.25; and, interpolating the R-band measurements of Galama et al., we derive R − I ∼ +2 on June 23. These colors, with the single exception of the R measurement of Galama et al. on June 21, are fully consistent with those of a late-type star such as LHS 2924. Very Large Array monitoring observations during the period June 19.5–25.7 show no source at the position of the OT with upper limits (1σ) of 0.05 mJy (at an observed frequency of 8.5 GHz), 0.05 mJy (at 5 GHz), and 0.1 mJy (at 1.4 GHz).”

1997 July 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green